A motor cyclist travels from A to B , at a speed of 30km/hr and returns to A at a speed of 50km/hr. Find the average speed and average velocity for whole journey.

Grade:11

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

Avg speed = total dist / total time taken
= 2d / ((d/30)+(d/50))
= 1500/40 km /hr = 37.5 km/hr

Avg velocity = ( net displacement/ total time)
= 0/ total time
= 0

Regards
Khimraj
4 years ago
verage required here is called the harmonic mean.

2/(1/30 + 1/50) = 37.5
...............................................................................................